OpenStudy (anonymous):

4t-7-t=7-2t-5 steps?

OpenStudy (anonymous):

t=9/5

OpenStudy (mathteacher1729):

There are a bunch of ways to solve it: 4t-7-t=7-2t-5 add 4t and -t on the left. 3t -7 = 7 - 2t - 5 add 7 to both sides 3t = -2t -5 add 2t to both sides 5t = -5 divide by 5 t = -1 There are other ways of course, but this is one.

OpenStudy (katrinakaif):

First combine the like terms: 4t - 7- t = 7 - 2t - 5 3t -7 = 2 - 2t Bring the "t"s all onto one side 3t + 2t -7 = 2 Bring the consants to one side 3t + 2t = 2+ 7 5t = 9 Divide by 5 t = 9/5

OpenStudy (mathteacher1729):

Hmm. Oh my goodness. My 2nd step had an error. :( Sorry. 3t -7 = 7 - 2t - 5 add 7 to both sides 3t = 14 -2t -5 add 2t to both sides

OpenStudy (anonymous):

KatrinaKaif how did u get 3t -7 = 2 - 2t?

OpenStudy (katrinakaif):

First simplifying the orginal equation in which the "t" on side could have been simplifyed (Collecting like terms)

OpenStudy (anonymous):

4t-7-t=7-2t-5 4t-t+2t =7-5+7 5t= 9 t=9/5